# Pack it in: Packing into Partially Filled Containers Through Contact

### v1

@@ II PROBLEM FORMULATION
We determine that an object $O$ at pose $T\in SE(3)$ is contained in the container if the volume of $O$, lies entirely within $C$ and its height is below the height threshold $t_{z}(T)<h_{max}$, where $t(T)\in\mathbb{R}^{3}$ denotes the translation component of a pose $T$, and $t_{z}(T)$ its vertical coordinate. The existing contents of the container are given by a set of $N$ rigid objects, $\mathcal{O}=\{O_{1},O_{2},\ldots,O_{N}\}$, where each object $O_{i}$ has known geometry and a 6D pose $T_{i}\in SE(3)$, specifying its position and orientation in the world frame, with $R_{i}\in SO(3)$ denoting the rotational component.
@@ III-A1 Residuals
  • End-Effector Force (EE FORCE): Magnitude of the 3D force vector at the end-effector tip: $\mathbf{f}_{ee}$.
@@ III-B2 Task Stages
Task Stage POS X POS Y POS Z UP EE FORCE RJL RJV OR
@@ IV-C Hardware Experiments
Method Success Force Failure Object Detached Exec Time (s) Dist Error (cm) EE Force (N)
@@
Force failure and object detached errors occur in a total of 20% of the failures for our PackItIn method. Force failures occurred when trying to displace an object away from the desired packing pose, when our trajectory optimization formulation instructed the robot to move in such a way that incurred large external forces due to object jamming. Notably, our method to penalise simulated EE forces appeared to reduce this effect during early testing, but was not able to completely eliminate this failure mode. The primary reason for object detached failure modes was when our assumption that the object was rigidly attached to the pump tip was violated significantly. Due to the minimalist pump we used, the object could rotate about the pump tip when it was used displace objects within the scene, if this rotation became too significant it could result in the object becoming removed from the pump tip.
@@ IV-E Hardware Failure Cases
where $T_{E}$ is the execution time of the trial, $N$ is the number of objects within the container, $t(T)^{a}_{t}$, $t(T)^{s}_{t}$ are the actual versus simulated positions of object $i$ at timestep $t$. $R_{\text{new},t}^{a}$, $R_{\text{new},t}^{s}$ were the actual versus simulated rotations of $O_{\text{new}}$ at timestep $t$. Finally $\mathbf{1}_{\cdot}$ represents a binary indicator function, returning 1 if the condition is true (OptiTrack perception failure).

### v2

@@ II PROBLEM FORMULATION
We determine that an object $O$ at pose $T\in SE(3)$ is contained in the container if the volume of $O$, lies entirely within $C$ and its height is below the height threshold $t_{z}(T)<h_{max}$, where $t(T)\in\mathbb{R}^{3}$ denotes the translation component of a pose $T$, and $t_{z}(T)$ its vertical coordinate. The existing contents of the container are given by a set of NN rigid objects, $\mathcal{O}=\{O_{1},O_{2},\ldots,O_{N}\}$, where each object $O_{i}$ has known geometry and a 6D pose $T_{i}\in SE(3)$, specifying its position and orientation in the world frame, with $R_{i}\in SO(3)$ denoting the rotational component.
@@ III-A1 Residuals
  • End-Effector Force ($E$ FORCE): Magnitude of the 3D force vector at the end-effector tip: $\mathbf{f}_{ee}$.
@@ III-B2 Task Stages
Task Stage POS X POS Y POS Z UP $E$ FORCE RJL RJV OR
@@ IV-C Hardware Experiments
Method Success Force Failure Object Detached Exec Time (s) Dist Error (cm) $E$ Force (N)
@@
Force failure and object detached errors occur in a total of 20% of the failures for our PackItIn method. Force failures occurred when trying to displace an object away from the desired packing pose, when our trajectory optimization formulation instructed the robot to move in such a way that incurred large external forces due to object jamming. Notably, our method to penalise simulated $E$ forces appeared to reduce this effect during early testing, but was not able to completely eliminate this failure mode. The primary reason for object detached failure modes was when our assumption that the object was rigidly attached to the pump tip was violated significantly. Due to the minimalist pump we used, the object could rotate about the pump tip when it was used displace objects within the scene, if this rotation became too significant it could result in the object becoming removed from the pump tip.
@@ IV-E Hardware Failure Cases
where $T_{E}$ is the execution time of the trial, NN is the number of objects within the container, $t(T)^{a}_{t}$, $t(T)^{s}_{t}$ are the actual versus simulated positions of object $i$ at timestep $t$. $R_{\text{new},t}^{a}$, $R_{\text{new},t}^{s}$ were the actual versus simulated rotations of $O_{\text{new}}$ at timestep $t$. Finally $\mathbf{1}_{\cdot}$ represents a binary indicator function, returning 1 if the condition is true (OptiTrack perception failure).